Hi Can I have some help with the following function?
Consider the following.
f(x) = -10x^2 + 6, g(x) = 3x - 7
(a) Find the function f compose g.

Answer by stanbon(75887)

f(x) = -10x^2 + 6, g(x) = 3x - 7
(a) Find the function f compose g.
-------------------------------
fog(x) = f[g(x)]
=f[3x-7]
= -10(3x-7)^2+6
= -10(9x^2-42x+49)+6
= -90x^2 + 420x -484
===========================
Cheers,
Stan H.